13.9 mol of helium is confined to a 19 L vessel at a pressure of 7 atm. Boltzmann’s constant is 1.38 × 10−23 J/K, and the universal gas constant is 8.31 J/K. Find the average kinetic energy of a helium molecule of mass 5.34 × 10−26 kg. Answer in units of J.
Answer #1

Solution:

Given;

no. of moles n=13.9

volume of the vessesl V=19L

Pressure P=7atm

Boltzmann's constant k=1.38\times 10^{-23}J/K

Gas constant R=8.31J/mol.K

Using the ideal gas equation, we can find the temperature of the vessel.

PV nRT

We have to use, the value of R=0.0821atm.L/mol.K .So,

  T=\frac{PV}{nR}=\frac{7\times 19}{13.9\times0.0821}=116.55K

The Average Kinetic energy of the helium molecule is given by,

K=\frac{3}{2}kT=\frac{3}{2}\times1.38\times10^{-23}\times116.55=2.41\times10^{-21}J

Thus, the Average kinetic energy of the helium molecule is 2.41\times10^{-21}J .